To expand operations, Aragon Consulting issued 1,600 shares of previously unissued common stock with a par value of $1. The price for the stock was $50 per share.


 1-a. Complete the table below, indicating the account, amount, and direction of the effect for the stock issuance.
Answer #1
Dear Student
Thank you for using homeworklib
Please find below the answer
Statementshowing Computations
Assets Liabilities
Cash = 1600*50 = 80,000 Common Stock = 1600*1 = 1600
Paid in capital = 1600*49 = 78,400
Transaction General Journal Debit Credit
                                            1.00 Cash dr               80,000.00
To Common Stock                  1,600.00
To Additional paid in capital               78,400.00